OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -5 f(1) = 20 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -5, -20, -65, -200, … -5, 20, -92, 372, … -5, -24, -92, -372, … -5, 20, -65, 200, …

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

i was thinking D

OpenStudy (skullpatrol):

why?

OpenStudy (anonymous):

if i knew i wouldnt be on here, lol, its a guess.

OpenStudy (anonymous):

ITs C probabily

OpenStudy (skullpatrol):

Is it obvious that the answer is either B or D?

OpenStudy (anonymous):

I dont even know, D stuck out to me

OpenStudy (anonymous):

Yeah I suck at math too...

OpenStudy (skullpatrol):

because the first two numbers are -5 an 20

OpenStudy (anonymous):

i guess so ... so is it D ?

OpenStudy (skullpatrol):

f(0) = -5 f(1) = 20

OpenStudy (skullpatrol):

So you need to compute f(2) using the given formula

OpenStudy (anonymous):

ok ...

OpenStudy (skullpatrol):

f(n) = -4•f(n -1) - 3•f(n - 2) f(2) = -4 * f(2-1) - 3 *f (2-2) f(2) = -4 * f(1) - 3 * f(0)

OpenStudy (anonymous):

ok..

OpenStudy (skullpatrol):

Is this making sense?

OpenStudy (anonymous):

not at all, looks like a different language, trying to explain it is pointless /:

OpenStudy (skullpatrol):

You are right, math is a different language :)

OpenStudy (skullpatrol):

A language of symbols.

OpenStudy (skullpatrol):

Take this formula $$f(2) = -4 * f(1) - 3 * f(0)$$ that I derived for you and substitute the values given from the question please.

OpenStudy (skullpatrol):

Remember: f(0) = -5 f(1) = 20

OpenStudy (anonymous):

i dont understand anything you just typed dude

OpenStudy (anonymous):

im putting D, i hope its the right answer

OpenStudy (skullpatrol):

Good guess.

OpenStudy (anonymous):

i hope it is

OpenStudy (skullpatrol):

Do you want to really try hard and understand this, if so I will explain it to you?

OpenStudy (anonymous):

it was right and no thanks

OpenStudy (skullpatrol):

@yeah_i_suck_at_math Do you still care?